Benfords Law claims that numbers chosen from very large data files tend to have 1 as the first nonzero digit disproportion
Answer #1

Given sample size n=345 Out of n, r=82 p=x/n = 82/345= 0.2377 To test Ho:P = 0.301 vs Hj: P<0.301 Test statistic Z= P-Po Po(1